# Explainability of AI Uncertainty: Application to Multiple Sclerosis Lesion Segmentation on MRI

### v1

@@ 3.2 Model
We chose nnU-Net, a widely adopted framework for medical image segmentation [27, 28], previously explored for lesion segmentation specifically [24, 44]. Based on the best practices for medical image segmentation, nnU-Net automatically configures pre-processing, training hyperparameters, and post-processing for a specific dataset. This approach ensures state-of-the-art results without extensive manual intervention, yet it allows for modifications on demand. We used the standard nnU-Net architecture operating on 3D full-resolution MP2RAGE scans.
@@ 3.2.1 Performance evaluation
Several measures chosen based on previous works were used to evaluate the model. These measures compare the predicted and ground-truth segmentation: (normalized) Dice similarity score (DSC and nDSC), lesion detection F1-score (LF1), precision (LPPV), and recall (LTPR). Analogously to DSC, nDSC measures the overall segmentation quality, however, it corrects for the difference in lesion loads across subjects, making their comparison fair [61]. For the lesion detection measures, the lesion regions were defined using the connected component analysis with 26 connectivity. True positive lesions were defined as lesions with a non-zero overlap with the ground truth. False positive lesions are predicted connected components without an overlap with the ground truth. False negative lesions are ground truth lesions without an overlap with the predicted lesion mask. The detailed metrics definition is described in A.
@@ 3.3.1 Deep ensemble
Deep ensembles for UQ have a wide adoption for medical imaging tasks given a strong uncertainty-robustness link and higher calibration compared to other methods, like temperature scaling, test time augmentation, Monte Carlo dropout, or stochastic weight averaging [8, 4]. Deep ensembles are a non-Bayesian UQ method based on training several models with different random seeds. Thus, each model in the ensemble has a distinct weight initialization, augmentations, training examples, shuffling, etc., providing convergence to different local minima. The resulting members’ weights are treated as sample estimations from the posterior distribution of model parameters. For each image voxel, uncertainty is computed as a ”spread” of the members’ predictions given the same example, i.e., using information theory measures (e.g., variance, mutual information, etc). We implemented UQ using a deep ensemble of $K=5$ models based on the nnU-Net framework architecture, training each model independently without using the bagging technique.
@@ 3.3.2 Lesion-scale uncertainty measures
Semantic segmentation task models output a tensor, associating class probabilities with each pixel or voxel of the input image. Such spatial nature allows uncertainty to be quantified at multiple levels: per voxel or ROI (e.g., focal lesions). While pixel-wise uncertainty provides detailed spatial information, instance-level uncertainty (computing a single uncertainty score for each predicted lesion) offers more clinically actionable insights. Computing a single uncertainty value per predicted lesion condenses predictive variability into compact, clinically meaningful information that aligns with how radiologists interpret and report findings, focusing on individual lesions rather than isolated voxels [13].
We quantify the uncertainty associated with each predicted CL using the lesion structural uncertainty (LSU) measure, which we previously proposed for WML segmentation [61]. LSU showed a slightly better ability to capture lesion false discovery errors. Additionally, having a simple interpretable definition and being bounded from 0 to 1, LSU allows for comparison across different patients. LSU quantifies the average disagreement between each ensemble member (k) lesion prediction, $L^{k},k\in\{0,2,\dots,K-1\}$, and the ensemble’s aggregated prediction $L$. Here, $L$ represents the final lesion segmentation obtained by averaging the probability maps from all $K$ ensemble members and applying a threshold to obtain a binary segmentation mask. LSU is defined as follows:
@@ 3.4.1 Feature engineering
Segmentation quality was computed as an adjusted intersection over the union (IoUadj[62] between the predicted lesion region ($L$) and the ground truth lesion mask. IoUadj is a suitable measure for medical image segmentation, where multiple ground truth instances can overlap with multiple predicted instances. IoUadj corrects the union computation, removing the ground truth part explained by other predicted instances. The IoUadj computation is detailed in B.
@@
The lesion location was quantified using the MNI atlas [63], separated into the right and left hemispheres. The location features were computed as the intersection between the lesion region and the MNI brain region divided by the lesion size in voxels. The location with respect to the cortex (i.e., cortex involvement) was computed as the intersection normalized by lesion size with the GM defined by SynthSeg [64]. In radiological practice, the unclear cortical involvement contributes to the aleatoric uncertainty in distinguishing between CL and WML that only touch the cortex (i.e., juxtocortical lesions).
@@
Radiomics features suggested by the PyRadiomics framework computes intensity and shape features for lesions and their surroundings [65]. Lesion surrounding was defined as a 4-voxel-wide region around the lesion, which was chosen in agreement with the medical doctors with an admitted degree of ambiguity. We used a morphological dilation with a $3\times 3\times 3$ voxels to obtain the lesion surrounding. PyRadiomics intensity features (first-order statistics) describe the distribution of voxel intensities within the ROI, such as mean, variance, skewness, and energy, providing insights into the overall brightness, variability, and histogram shape. The texture features (second-order statistics) analyze the spatial relationship and patterns of voxel intensities at specific distances and directions. We focused on the texture features derived from the gray-level co-occurrence matrix (GLCM) with a distance of 1 voxel in all 13 3D directions, quantifying contrast, correlation, energy, and homogeneity. The rest of the texture features are omitted given the low interpretability and to limit the amount of features below the number of predicted lesions. The shape features capture the 3D geometry of the ROI and focus on the structural and morphological aspects, e.g., volume, surface area, fraction (compactness, sphericity), and longest or shortest axes, among others.
@@
The degree of novelty describes the difference between a lesion and lesions in the Train set. To measure the lesion novelty, we computed the distance between the lesion and the training set using the engineered lesion and perilesional features. For dimensionality and noise reduction, probabilistic principal component analysis (PPCA) [66] was applied to the features before computing the distance. PPCA provides a maximum likelihood estimation approach to dimensionality reduction while accounting for noise in the data. We used two measures to compute the distance: Mahalanobis distance and non-zero distance to the closest example in the training set.
@@
Patient-wise measures are also used to examine how global features are related to the lesion uncertainty. First, we added PyRadiomics intensity statistics for the whole predicted lesion region. Second, to study the relationship between uncertainty and image quality, we relied on MRIQC [67], characterizing various aspects of MR image quality such as signal-to-noise ratio, motion artifacts, ghosting, and entropy in the whole brain and the specific brain tissues. We selected 13 metrics applicable for structural skull-stripped MRI scans (the selection procedure is detailed in the Supplementary materials). Finally, we used patient demographic and clinical information, such as sex, age, and expanded disability status scale (EDSS).
@@ 3.4.3 Uncertainty regression model
Linear model. We chose Elastic-Net [68] model, which combines L1 and L2 regularization, providing feature selection and penalizing for multi-collinearity, respectively. Given the higher prevalence of low-uncertainty lesions and a greater interest in explaining high-uncertainty lesions, we implemented an increased penalty for high-uncertainty examples, i.e., sample weighing based on uncertainty in the loss function and the quality metrics. The model selection procedure was performed using grid search, optimizing the weighted $R^{2}$ score obtained on 5-fold cross-validation. The hyperparameters tuned during the model selection control the regularization strength (9 values from $10^{-4}$ to $10^{2}$ evenly spaced on the log scale), the L1 to L2 regularization strength ratio (9 values evenly spaced between 0 and 1), and the tolerance for the optimization ($10^{-3}$, $10^{-4}$, $10^{-5}$). To obtain robust feature importance values, the grid search procedure was repeated 10 times with different random seed initializations, controlling the data-splitting and Elastic-Net convergence. Additionally, to study the relationship between the prediction quality and uncertainty, we used the ordinary least squares regression model. The analysis was repeated 10 times with the random seed change affecting the cross-validation data-splitting.
Random forests. Random forests [69] is an ensemble learning method that builds multiple decision trees during training, using random subsets of data and features, and combines their outputs (via averaging for regression or majority voting for classification) to make predictions. Feature importances in random forests are computed by evaluating how much each feature reduces the impurity (in our regression case, variance) across all trees in the forest. The model selection procedure was performed using grid search, optimizing the $R^{2}$ score obtained on 5-fold cross-validation. The search grid was defined by varying the number of estimators (20, 50, 100), the maximum depth of the trees (none, 5, 10), the minimum number of samples required to split an internal node or to be at a leaf node (2, 5), and the number of considered features for splitting (N, N²). Analogously to the linear model, the grid search was repeated 10 times with different random seed initialization.
@@ 3.5.1 Quality of fit
The quality of fit as a proxy to the quality of explanations was computed for each of 10 runs, either during the 5-fold cross-validation or over a dataset (Train, Test-in, Test-out) for the best selected model. We reported the average and standard deviation across 10 runs with different random seeds.
@@ 4.1 Deep learning segmentation model performance
Table 3 reports CL segmentation and detection performance of the trained nnU-Net model. Judging by the Dice similarity score (DSC) and detection F1, the performance on Test-out was significantly worse than on Test-in. This confirms the suitability of Test-out for investigating the quality of uncertainty out-of-domain, since it enables the evaluation of uncertainty explanations and sanity in the scenario of model failure. The total number of predicted lesions is also reported as the number of examples used to fit the explainer model.
Table 3: Mean and standard deviation of the segmentation (DSC, nDSC) and lesion detection quality metrics (LF1, LPPV, LTPR) across Train, Test-in, and Test-out datasets, as well as the number of predicted CL (# CLs).
Dataset DSC nDSC LF1 LPPV LTPR # CLs
Train 0.908 $\pm$ 0.082 0.740 $\pm$ 0.289 0.957 $\pm$ 0.078 0.991 $\pm$ 0.030 0.934 $\pm$ 0.120 753
Test-in 0.617 $\pm$ 0.332 0.494 $\pm$ 0.380 0.728 $\pm$ 0.300 0.759 $\pm$ 0.333 0.778 $\pm$ 0.287 284
Test-out 0.318 $\pm$ 0.373 0.284 $\pm$ 0.372 0.399 $\pm$ 0.396 0.529 $\pm$ 0.464 0.377 $\pm$ 0.406 273
@@ 4.2 Performance in explaining uncertainty
For the linear model, using all features, the $R^{2}$ was highest when fitted on the Train set; it decreased by 20% on Test-in and by an additional 33% on Test-out. The MAE increased by 50% from Train to Test-in and by 17% from Test-in to Test-out. The standard deviation of $R^{2}$ on the Test-out set is high due to an outlier appearing under a specific random seed. However, this outlier did not affect the feature importance analysis (see Supplementary materials).
The quality of fit, measured by $R^{2}$, more than doubled when all features were used compared to using only lesion segmentation quality to predict uncertainty. This suggests that the designed features captured more variability in uncertainty than lesion segmentation quality (IoUadj) alone. Fitting the explainer on all features, but the predicted quality led to up to an 11% decrease in the $R^{2}$ compared to using all features on Test-in and Test-out. This indicates that the proposed features could not completely replace the information about the segmentation error. When patient scale information was omitted, $R^{2}$ improved by 2% on Train and 3% on Test-out, but decreased by 2% on Test-in.
@@
The random forest quality of fit is shown in Table 5. MAE values of random forest models are comparable with the values for the linear models for Train and Test-in. The $R^{2}$ is lower compared to the linear model for each feature set: up to 15% lower on Train and Test-in, up to 60% lower on Test-out. For all datasets, removing either patient-scale information or the prediction quality led to lower $R^{2}$ values compared to using all features; however, the difference was marginally small for Train and Test-in. Thus, further, we reserved the linear model for explanations in the setting with all features. Despite slightly lower performance, the standard deviation of the cross-validation scores across different seeds is lower for the random forest model compared to linear models.
@@
We analyze the linear model prediction quality on sets different from the fitting sets and evaluate the quality of explanations under the domain shift. The results are shown in Figure 4. A drop in performance caused by the domain shift was observed for all fitting settings, except for one. When IoUadj was used to predict uncertainty using Test-out for fitting, the $R^{2}$ was higher on Train and Test-in compared to Test-out. This should indicate a weaker relationship between uncertainty and error on the out-of-domain set, compared to in-domain sets. The drop in performance tended to be worsened by the out-of-domain setting. Thus, explanations on the Train and Test-in sets should be more similar to each other than to Test-out.

### v2

@@ 3.2 Model
We chose nnU-Net, a widely adopted framework for medical image segmentation  (Isensee et al., 2021; González et al., 2023), previously explored for lesion segmentation specifically  (Rondinella et al., 2025; Molchanova et al., 2025). Based on the best practices for medical image segmentation, nnU-Net automatically configures pre-processing, training hyperparameters, and post-processing for a specific dataset. This approach ensures state-of-the-art results without extensive manual intervention, yet it allows for modifications on demand. We used the standard nnU-Net architecture operating on 3D full-resolution MP2RAGE scans.
@@ 3.2.1 Performance evaluation
Several measures chosen based on previous works were used to evaluate the model. These measures compare the predicted and ground-truth segmentation: (normalized) Dice similarity score (DSC and nDSC), lesion detection F1-score (LF1), precision (LPPV), and recall (LTPR). Analogously to DSC, nDSC measures the overall segmentation quality, however, it corrects for the difference in lesion loads across subjects, making their comparison fair  (Raina et al., 2023). For the lesion detection measures, the lesion regions were defined using the connected component analysis with 26 connectivity. True positive lesions were defined as lesions with a non-zero overlap with the ground truth. False positive lesions are predicted connected components without an overlap with the ground truth. False negative lesions are ground truth lesions without an overlap with the predicted lesion mask. The detailed metrics definition is described in A. Mann-Whitney U-tests with Benjamini-Hochberg false discovery rate correction ($\alpha=0.05$) were conducted to check for significant performance changes across test sets.
@@ 3.3.1 Deep ensemble
Deep ensembles for UQ have a wide adoption for medical imaging tasks given a strong uncertainty-robustness link and higher calibration compared to other methods, like temperature scaling, test time augmentation, Monte Carlo dropout, or stochastic weight averaging  (Gawlikowski et al., 2023; Lambert et al., 2024). Deep ensembles are a non-Bayesian UQ method based on training several models with different random seeds. Thus, each model in the ensemble has a distinct weight initialization, augmentations, training examples, shuffling, etc., providing convergence to different local minima. The resulting member weights are treated as sample estimations from the posterior distribution of model parameters. For each image voxel, uncertainty is computed as a “spread” of the member predictions given the same example, i.e., using information theory measures (e.g., variance, mutual information, etc). We implemented UQ using a deep ensemble of $K=5$ models based on the nnU-Net framework architecture, training each model independently without using the bagging technique.
@@ 3.3.2 Lesion-scale uncertainty measures
Semantic segmentation task models output a tensor, associating class probabilities with each pixel or voxel of the input image. Such spatial nature allows uncertainty to be quantified at multiple levels: per voxel or ROI (e.g., focal lesions). While pixel-wise uncertainty provides detailed spatial information, instance-level uncertainty (computing a single uncertainty score for each predicted lesion) offers more clinically actionable insights. Computing a single uncertainty value per predicted lesion condenses predictive variability into compact, clinically meaningful information that aligns with how radiologists interpret and report findings, focusing on individual lesions rather than isolated voxels  (Huet-Dastarac et al., 2024).
We quantify the uncertainty associated with each predicted CL using the lesion structural uncertainty (LSU) measure, which we previously proposed for WML segmentation  (Raina et al., 2023). LSU showed a slightly better ability to capture lesion false discovery errors. Additionally, having a simple, interpretable definition and being bounded from 0 to 1, LSU allows for comparison across different patients. LSU quantifies the average disagreement between each ensemble member (k) lesion prediction, $L^{k},k\in\{0,2,\dots,K-1\}$, and the ensemble’s aggregated prediction $L$. Here, $L$ represents the final lesion segmentation obtained by averaging the probability maps from all $K$ ensemble members and applying a threshold to obtain a binary segmentation mask. LSU is defined as follows:
@@
We employed a standard evaluation strategy for structure-wise uncertainty measures by computing Spearman’s rank correlation between predicted uncertainty and lesion-wise prediction quality (Roy et al., 2019; Lambert et al., 2022). This approach assesses the extent to which uncertainty correlates with prediction error, based on the premise that higher values from a reliable uncertainty measure should correspond to a greater likelihood of incorrect predictions. Details on the lesion prediction quality metric are provided in the following section.
@@ 3.4.1 Feature engineering
Segmentation quality was computed as an adjusted intersection over the union (IoUadj(Rottmann et al., 2020) between the predicted lesion region ($L$) and the ground truth lesion mask. IoUadj is a suitable measure for medical image segmentation, where multiple ground truth instances can overlap with multiple predicted instances. IoUadj corrects the union computation, removing the ground truth part explained by other predicted instances. The IoUadj computation is detailed in B.
@@
The lesion location was quantified using the MNI atlas (Grabner et al., 2006), separated into the right and left hemispheres. The location features were computed as the intersection between the lesion region and the MNI brain region, divided by the lesion size in voxels. The location with respect to the cortex (i.e., cortex involvement) was computed as the intersection normalized by lesion size with the GM defined by SynthSeg  (Billot et al., 2023). In radiological practice, the unclear cortical involvement contributes to the aleatoric uncertainty in distinguishing between CL and WML that only touch the cortex (i.e., juxtocortical lesions).
@@
Radiomics defined by the PyRadiomics framework computes intensity and shape features for lesions and their surroundings  (van Griethuysen et al., 2017). Perilesional region (i.e., lesion surrounding) was defined as a 4-voxel-wide region around the lesion, which was chosen in agreement with the medical doctors with an admitted degree of ambiguity. We used a morphological dilation with a $3\times 3\times 3$ voxels to obtain the lesion surrounding. PyRadiomics intensity features (first-order statistics) describe the distribution of voxel intensities within the ROI, such as mean, variance, skewness, and energy, providing insights into the overall brightness, variability, and histogram shape. The texture features (second-order statistics) analyze the spatial relationship and patterns of voxel intensities at specific distances and directions. We focused on the texture features derived from the gray-level co-occurrence matrix (GLCM) with a distance of 1 voxel in all 13 3D directions, quantifying contrast, correlation, energy, and homogeneity. The rest of the texture features are omitted, given the low interpretability and limited number of features below the number of predicted lesions. The shape features capture the 3D geometry of the ROI and focus on the structural and morphological aspects, e.g., volume, surface area, fraction (compactness, sphericity), and longest or shortest axes, among others.
@@
The degree of novelty describes the difference between a lesion and lesions in the Train set. To measure the lesion novelty, we computed the distance between the lesion and the training set using the engineered lesion and perilesional features. For dimensionality and noise reduction, probabilistic principal component analysis (PPCA)  (Bishop, 2006) was applied to the features before computing the distance. PPCA provides a maximum likelihood estimation approach to dimensionality reduction while accounting for noise in the data. We used two measures to compute the distance: Mahalanobis distance and non-zero distance to the closest example in the training set.
@@
Patient-wise measures are also used to examine how global features are related to the lesion uncertainty. First, we added PyRadiomics intensity statistics for the whole predicted lesion region. Second, to study the relationship between uncertainty and image quality, we relied on MRIQC (Esteban et al., 2017), characterizing various aspects of MR image quality such as signal-to-noise ratio, motion artifacts, ghosting, and entropy in the whole brain and the specific brain tissues. We selected 13 metrics applicable for structural skull-stripped MRI scans (the selection procedure is detailed in the Supplementary materials). Finally, we used patient demographic and clinical information, such as sex, age, and expanded disability status scale (EDSS).
@@ 3.4.3 Uncertainty regression model
Linear model. We chose Elastic-Net  (Zou and Hastie, 2005) model, which combines L1 and L2 regularization, providing feature selection and penalizing for multi-collinearity, respectively. Given the higher prevalence of low-uncertainty lesions and a greater interest in explaining high-uncertainty lesions, we implemented an increased penalty for high-uncertainty examples, i.e., sample weighing based on uncertainty in the loss function and the quality metrics. The model selection procedure was performed using grid search, optimizing the weighted $R^{2}$ score obtained on 5-fold cross-validation. The hyperparameters tuned during the model selection control the regularization strength (9 values from $10^{-4}$ to $10^{2}$ evenly spaced on the log scale), the L1 to L2 regularization strength ratio (9 values evenly spaced between 0 and 1), and the tolerance for the optimization ($10^{-3}$, $10^{-4}$, $10^{-5}$). To obtain robust feature importance values, the grid search procedure was repeated 10 times with different random seed initializations, controlling the data-splitting and Elastic-Net convergence. Additionally, to study the relationship between the prediction quality and uncertainty, we used the ordinary least squares regression model. The analysis was repeated 10 times with the random seed change affecting the cross-validation data splitting.
Random forests. Random forests  (Breiman, 2001) is an ensemble learning method that builds multiple decision trees during training, using random subsets of data and features, and combines their outputs (via averaging for regression or majority voting for classification) to make predictions. Feature importances in random forests are computed by evaluating how much each feature reduces the impurity (in our regression case, variance) across all trees in the forest. The model selection procedure was performed using grid search, optimizing the $R^{2}$ score obtained on 5-fold cross-validation. The search grid was defined by varying the number of estimators (20, 50, 100), the maximum depth of the trees (none, 5, 10), the minimum number of samples required to split an internal node or to be at a leaf node (2, 5), and the number of considered features for splitting (N, N²). Analogously to the linear model, the grid search was repeated 10 times with different random seed initializations.
@@ 3.5.1 Quality of fit
The quality of fit, used as a proxy for explanation quality, was computed over 10 independent runs—either within the 5-fold cross-validation setting or across dataset splits (Train, Test-in, Test-out) for the best-selected model. We report the mean and standard deviation across runs initialized with different random seeds. We conducted non-parametric statistical tests to assess the significance of the fitting quality differences across models (paired Wilcoxon signed-rank test) and datasets (Mann-Whitney U-test).
@@ 4.1 Deep learning segmentation model performance
Table 3 reports CL segmentation and detection performance of the trained nnU-Net model. The segmentation and lesion detection performance are significantly worse on Test-out compared to Train ($\textit{P}<0.01$) and Test-in ($\textit{P}<0.05$ for Precision and $\textit{P}<0.01$ for other metrics). This enables the evaluation of uncertainty explanations and sanity under a strong domain shift leading to model failure.
Table 3 also reports the Spearman’s correlation between uncertainty and error per dataset (Unc.-error $\rho$). LSU computed for Train and Test-in set passes the uncertainty sanity check, correlating higher than 0.5; on the Test-out $\rho$ to 0.35, signaling a weaker relationship between uncertainty and error out-of-domain.
The number of predicted CLs, i.e., the number of examples used to fit the explainer model, per dataset is 753 lesions for Train, 284 for Test-in, and 273 for Test-out.
All P-values from the statistical tests are provided in the Supplementary Materials.
Table 3: Mean and standard deviation of the segmentation (DSC, nDSC) and lesion detection quality metrics (LF1, LPPV, LTPR) across Train, Test-in, and Test-out datasets, as well as the number of predicted CL (# CLs) and correlation between LSU and IoUadj (Unc.-error $\rho$).
Dataset DSC nDSC LF1 LPPV LTPR # CLs Unc.-error $\mathbf{\rho}$
Train 0.908 $\pm$ 0.082 0.740 $\pm$ 0.289 0.957 $\pm$ 0.078 0.991 $\pm$ 0.030 0.934 $\pm$ 0.120 753 0.543
Test-in 0.617 $\pm$ 0.332 0.494 $\pm$ 0.380 0.728 $\pm$ 0.300 0.759 $\pm$ 0.333 0.778 $\pm$ 0.287 284 0.545
Test-out 0.318 $\pm$ 0.373 0.284 $\pm$ 0.372 0.399 $\pm$ 0.396 0.529 $\pm$ 0.464 0.377 $\pm$ 0.406 273 0.353
@@ 4.2 Performance in explaining uncertainty
For the linear model, using all features, the $R^{2}$ was highest when fitted on the Train set; it decreased by 20% on Test-in and by an additional 33% on Test-out. The MAE increased by 50% from Train to Test-in and by 17% from Test-in to Test-out. These drops in the performance across the domains are statistically significant according Mann-Whitney U-test with $\textit{P}<0.001$. The standard deviation of $R^{2}$ on the Test-out set is high due to an outlier appearing under a specific random seed. However, this outlier did not affect the feature importance analysis (see Supplementary materials).
The quality of fit, measured by $R^{2}$, more than doubled when all features were used compared to using only lesion segmentation quality to predict uncertainty. The change is statistically significant according to the paired Wilcoxon tests ($\textit{P}<0.01$) for all datasets. This suggests that the designed features captured more variability in uncertainty than lesion segmentation quality (IoUadj) alone. Fitting the explainer on all features, but the predicted quality, led to up to an 11% decrease in the $R^{2}$ compared to using all features on Test-in and Test-out ($\textit{P}<0.01$), but no significant change on Train. This indicates that the proposed features could not completely replace the information about the segmentation error. When patient scale information was omitted, $R^{2}$ improved by 2% on Train ($\textit{P}=0.006$) and 3% on Test-out ($\textit{P}=0.431$), but decreased by 2% on Test-in ($\textit{P}=0.232$).
@@
The random forest quality of fit is shown in Table 5. The MAE values of random forest models are comparable to those of the linear models for Train and Test-in. The differences in MAE on Train and Test-in are not statistically significant according to the Wilcoxon rank test ($\textit{P}>0.1$) for all settings, but Train with only lesion-scale features ($\textit{P}=0.006$). The $R^{2}$ is lower than the linear model for each feature set: up to 15% lower on Train and Test-in ($\textit{P}<0.01$), up to 60% lower on Test-out ($\textit{P}<0.1$). For all datasets, removing either patient-scale information or the prediction quality led to lower $R^{2}$ values compared to using all features. However, the difference in $R^{2}$ is statistically significant only when removing IoUadj and fitting on Test-in and when removing patient information and fitting on Test-out ($\textit{P}<0.01$). Despite slightly lower performance, the standard deviation of the cross-validation scores across different seeds is lower for the random forest model than for linear models.
@@
Given these results, we chose the ElasticNet model with all features as the main explanatory model. We analyzed this model prediction quality on sets different from the fitting sets and evaluated the quality of explanations under the domain shift. The results are shown in Figure 4. A drop in performance caused by the domain shift was observed for all fitting settings, except for one. When IoUadj was used to predict uncertainty using Test-out for fitting, the $R^{2}$ was higher on Train and Test-in than on Test-out, indicating a weaker relationship between uncertainty and error on the out-of-domain set, compared to in-domain sets. Overall, the drop in performance tended to be worsened by the out-of-domain setting, and the differences in the performance are statistically significant with $\textit{P}<0.01$ according to the Mann-Whitney U-test. Thus, explanations on the Train and Test-in sets should be more similar to each other than to Test-out.
The exact P-values for all statistical tests are provided in the Supplementary material.